how do you convert 4 1/8 pounds to ounces.
Answer by Theo(2978)

convert 4 + 1/8 pounds to ounces.

there are 16 ounces to the pound.

the ratio of ounces to pounds is 16/1.

multiply your pounds by 16/1 to get ounces.

4 + 1/8 pounds * 16 = 16*4 + 16*(1/8) = 64 + 2 = 66 ounces

to convert this back to pounds, divide by 16 to get:

66 / 16 = 4.125 pounds.

that's equal to 4 pounds plus .125 pounds.

multiply .125 * 8/8 to get 1/8.

your answer of 66 ounces is back to 4 and 1/8 pounds which confirms your answer is good.

your answer is 66 ounces.